OpenStudy (anonymous):

differentiate y=3sinx + 4tanx

OpenStudy (anonymous):

piece by piece the derivative of sine is ?

OpenStudy (anonymous):

cosx right?

OpenStudy (anonymous):

@satellite73

zepdrix (zepdrix):

yes! Do you know the derivative of tanx? If not, you can write it as \(\Large\rm \frac{\sin x}{\cos x}\) and apply the quotient rule :)

OpenStudy (anonymous):

\[\sec ^{2}\]

OpenStudy (anonymous):

and from the sec2. how can i make it in simplest way?

zepdrix (zepdrix):

That would be the simplest form for the derivative.

OpenStudy (anonymous):

\[\sec ^{2} is \it equal \to 1+\tan ^{2} ?\]

zepdrix (zepdrix):

\[\Large\rm y=3\sin x+4 \tan x\]\[\Large\rm y'=3\cos x+4\sec^2x\]You really can't make it simpler than this :d Unless you really really want your answer in terms of sines and cosines.

OpenStudy (anonymous):

owhhh.. so it's ok if i just left it like that? thanks. really appreciate ur help..

zepdrix (zepdrix):

ya c: